OpenStudy (anonymous):

What is the value of x? 1/4x - 8 + 2/4x=19

OpenStudy (anonymous):

@AlexandervonHumboldt2

OpenStudy (anonymous):

plz help

OpenStudy (anonymous):

I will try

OpenStudy (anonymous):

Medal to the best answer and fan

OpenStudy (anonymous):

kay

OpenStudy (anonymous):

x in (-oo:+oo) (1/4)*x+(2/4)*x-8 = 19 // - 19 (1/4)*x+(2/4)*x-19-8 = 0 3/4*x-27 = 0 // + 27 3/4*x = 27 // : 3/4 x = 27/3/4 x = 36 x = 36

OpenStudy (misty1212):

it is not exactly clear what the equation is

OpenStudy (anonymous):

tell when you medal me

OpenStudy (anonymous):

yeah huh

OpenStudy (anonymous):

and fan me

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

your welcome, did you fan me and give me a medal

OpenStudy (anonymous):

YES